A paddle boat can move at a speed of 20 km/h in still water. The boat is paddled 14 km downstream in a river in the same time it takes to go 7 km upstream. What is the speed of the river?
Answer by mananth(16949)

Boat speed = 20 km/h
current speed = x km/h

speed against current 20 - x
speed with current 20 + x

Distance against current 7 km
Distance with current 14 km

t=d/r time against = time with current
7/(20-x )=14/(20+x)
7(20+x)=14(20-x)
140+7x =280-14 x
14x+7x=280-140
21x=140
/21
x = 6.67 km/h current speed